Consider the titration of 40 mL of 0.114 MNaOH with 0.0790MHCl. Calculate the pH after the addition of each of the following volumes of acid

a) 7.4ml (b) 40ml  (c) .12L
Answer #1

a)

millimoles of NaOH = 40 x 0.114 = 4.56

millimoles of HCl = 7.4 x 0.0790 = 0.5846

[OH-] = 4.56 - 0.5846 / 40 + 7.4

         = 0.0838 M

pOH = -log [OH-] = -log (0.08387)

        = 1.08

pH = 12.92

b)

mmoles of HCl = 40 x 0.0790 = 3.16

[OH-] = 4.56 - 3.16 / 40 + 40 = 0.0175 M

pOH = 1.76

pH = 12.24

c)

mmoles of HCl = 0.12 x 1000 x 0.0790 = 9.48

[H+] = 9.48 - 4.56 / 120 + 40 = 0.03075 M

pH = 1.51